### v1

# sample.bib

Abstract
[…]
Enlightenment English philosophy characterizes God as progressively more impersonal, with “Nature” and “Law” acting as substitutions for the imminent Creator [darwin1859origin, DeclarationofIndependence:atranscription_2024]:
But with regard to the material world, we can at least go so far as this - we can perceive that events are brought about not by insulated interpositions of Divine power, exerted in each particular case, but by the establishment of general laws.
—Anglican William Whewell quoted on the first page of Darwin’s Origin of Species
…and equal station to which the Laws of Nature and of Nature’s God entitle them…
—Thomas Jefferson in the Declaration of Independence, a deist and nominally a member of the Church of England
and Wolfram himself writes [Howwegotherewolfram]:
As the centuries went by, the idea of “natural laws” sharpened, sometimes with an almost computational feel. “God can only run the world by natural laws”, or “The universe is the thoughts of God actualized”
So—perhaps influenced by the secular milieu of mid-to-late 20th century science—he develops this idea further: God did not give rules, God
is rules! And by doing so, it would seem that he completes the removal of anthropocentrism in creation and further divides what Heisenberg calls the Cartesian partition (“God-World-I”) [Heisenberg]. But this assessment is not correct, because the dependence of mathematical truths on the experience of “observers like us”(sic) is a major conclusion of Wolfram’s work [Somehistoricalbackgroundwolfram]:
The intellectual foundations and justification are different now. But in a sense our view of mathematics has come full circle. And we can now see that mathematics is in fact deeply connected to the physical world and our particular perception of it. And we as humans can do what we call mathematics for basically the same reason that we as humans manage to parse the physical world to the point where we can do science about it.
How far does “our particular perception” allow us to speak on fundamental theories of physics and metaphysics? According to Wolfram, the underlying substrate of reality is the “entangled limit of all possible computations” called “the ruliad”. This conjecture is either as abstract as the unknowable God and unfalsifiable, or so vacuous one might as well say “the universe does things”, or some superposition of the two. Tegmark, on the other hand, asserts reality is all “mathematical structure”—a form of radical platonism that aligns with his acceptance of the Many-Worlds interpretation of quantum mechanics [Tegmark_1998]. Yes, a culmination of phenomena including quantum mechanics, the modern computer, and lack of alternative dogma have enabled what I might describe as gnostic atomism. The following paragraph will explicate a brief history of similar ideas, although a more in-depth history is found in Heisenberg’s book on the philosophy of physics [Heisenberg].
Democritus is the most well-known earliest proponent of the theory of the atom around 400 BC (who, like Wolfram and Tegmark, was controversial) and his theory is roughly aligned with modern atomic theories of today [AncientAtomism]. In some ways modern theories supersede those of the past, but in other ways they are no different. For example, Aristotelian and Asharite philosophers described reality in terms of atoms and “accidents” [ArabicandIslamicNaturalPhilosophy], Democritus described atoms and “hooks” [Democritus] which are both similar to the idea of vertices and edges on Wolfram’s hypergraphs. It is unsurprising that Wolfram’s use of natural philosophy and transcendent ruliad bear similarity to Islamic ideas. Abu al-Hudhayl—regarded as the founder of kalām atomism—posited that atoms were indiscernible from one another in size or shape, similar to emes, Wolfram’s fundamental atom of computation. Emes are given unique identifiers, which is comparable to Leibniz’s Monads where “each Monad must be different from every other” and he describes them as “incorporeal automata” [Leibniz_1714]. While atomism of the past relied on crude arguments using primitive technology such as grain mills [Dhanani_2015] or steam engines, Wolfram and Tegmark use the digital computer to animate their ideas. The advantage of the modern computer for Wolfram’s theory is that it allows for large combinatorial statistics experiments to be run, which form the basis for reproducing all natural laws of 20th century physics that physicists care about. More recent developments for discrete models of our universe come from Rafael Sorkin’s causal set theory [causalset] and Ed Fredkin’s SALT cellular automaton model. Before them, John Von Neumann explored self-replicating cellular automata with the goal of designing a machine whose complexity could grow automatically akin to biological organisms under natural selection [neumannautomata].
In this work I will undertake a similar project to philosopher of science Nancy Cartwright who aligns with me in the “Aristotelian belief in the richness and variety of the concrete and particular” [cartwrightlies] and will reframe some of the challenges put forward by Aristotle in Metaphysics [metaphysicsaristotle]. It was Aristotle who criticized both Plato and fourth-century Pythagoreans for constructing natural bodies possessing weight from indivisible mathematical abstractions [Pythagoreanism]. An error of the “so-called” Pythagoreans is described in [platopythag]:
The “so-called” Pythagoreans adapt the immediate facts to fit the explanation in an ad hoc manner. The example given involves the bodies of the heavens: nine bodies can be perceived by the senses, but, since the Pythagoreans assume the number 10 as the perfect number, and since all things are number, there must be ten heavenly bodies. This example also reveals Aristotle’s second substantial criticism of the philosophical system of the “so-called” Pythagoreans: they hastily and carelessly compare things in order to secure relationships between their first principles and observed phenomena.
We begin by discussing the current assertions given by Wolfram himself. In the model we assume observers like us have [ObserverTheory—StephenWolframWritings_2023]
  1. 1.
    Persistence in time
  2. 2.
    Persistence in space
  3. 3.
    “Independence” from the universe
  4. 4.
    Stability of discrete symbolic concepts
  5. 5.
    Independence from the hyperruliad
Tegmark shares the first four beliefs. Some assertions of Wolfram regarding these assumptions are as follows:
  • But why should that summary (the perception of our universe) have any coherence? Basically it’s because we impose coherence through our definition of how observers like us work.
  • But actually there’s more than just our assumption of persistence in time. For example, we also have an assumption of persistence in space: we assume that—at least on reasonably short timescales—we’re consistently “observing the universe from the same place”, and not, say, “continually darting around”.
  • Because in fact observers that are even vaguely like us are in effect deeply constrained in what rules they can attribute to the universe.
  • But we make the implicit assumption that, no, the universe at least as far as we perceive it - is a more organized and consistent place. And indeed it’s that assumption that makes it feasible for us to operate as observers like us at all, and to even imagine that we can usefully reduce the complexity of the world to something that “fits in our finite minds”.
  • We assume that we can take all the complexity of the world and represent at least the parts of it that we care about in terms of discrete symbolic concepts, of the kind that appear in human (or computational) language.
  • These assertions seem very general, and in some ways almost self-evident-at least as they apply to us.
In using the plural “us” and “our” consistently, we can assume Stephen is assuming a common agreement amongst readers, or is employing nosism. Either way, the presumption that many humans have a common sampling of the ruliad is left unexplained, along with why we are and remain localized in space and time. It is unclear how observers are delineated in the ruliad and how this is compatible with the purported independence from the universe—the idea of “observers like us” is simply a brute fact for Wolfram. These proposed restrictions on observers bring to mind the fine-tuning argument for intelligent design: as observers, we happen to be just the right kind to be able to name the underlying structure of reality itself. To quote Wolfram, “There has to be just enough equivalencing” of information by human observers “and not too much.” He admits circularity in his reasoning himself:
Well, the point is that an observer can equivalence those successive patterns of emes, so that what they observe is persistent. And, yes, this is at least on the face of it circular. And ultimately to identify what parts of the ruliad might be “persistent enough to be observers”, we’ll have to ground this circularity in some kind of further assumption.
Figure 1: An interpretation of the metaphysics given by Wolfram’s ruliad and Tegmark’s MUH that exhibits circularity.
So Wolfram’s current universe is dyadic—it consists of the ruliad and observers. If conscious observers were emes, these emes would still follow privileged rules which uphold the aforementioned assumptions regarding the observer. For example, it could be that observers are time-and-space-localized compression emes which sample the ruliad. But Wolfram (and Tegmark) are notably unsatisfied with a privileged rule which runs the universe:
At the outset, we might have imagined that the end point of our project would be the identification of some particular rule of which we could say “This is the rule for the universe”. But of course then we’d be faced with the question: “Why that rule, and not another?” And perhaps we would imagine just having to say “That’s something that you have to go beyond science to answer”.
Wolfram and Tegmark avoid this challenge like the “so-called” Pythagoreans in the aforementioned example by claiming that instead of one rule, all rules are run. Wolfram forces the mathematical physics jargon “in the entangled limit”. I suspect by invoking terms used often in quantum mechanics and calculus, the reader should naturally intuit what he means.
Let’s turn to more technical definitions of the observer as given by Gorard [gorard1]:
[…]
Definition 24 An “observer” is any persistent structure within a multiway system that perceives a single, definitive evolution history.
Definition 37 An “observer” in a multiway system is any ordered sequence of non-intersecting branchlike hypersurfaces $\Sigma t$ that covers the entire multiway graph, with the ordering defined by some universal time function…
Critical to the model is the abstract rewriting system [gorard2]:
Definition 5 An “abstract rewriting system” is a set, denoted A (where each element of A is known as an “object”), equipped with some binary relation, denoted $\rightarrow$, known as the “rewrite relation”.
How do we corroborate Stephen’s assertions with Gorard’s definitions and our conscious experience? From the definitions the model of reality is relations between vertices on hypergraphs. A binary relation is defined as a set of ordered pairs of elements from two sets. So from the eternally existing atoms of space spring forth such relations. Already the model seems vacuous because which rules are applied when presumably must evolve according to some hyperrule in the enigmatic and so-called “hyperruliad”—an infinite regress problem arises. Regarding the hyperruliad existing, Wolfram writes the following [Wolfram_ruliad]:
Yes, we could imagine some other entity that’s embedded within the hyperruliad, and perceives what it considers to be the universe to operate hypercomputationally. But in a statement that’s in a sense more “about us” than “about the universe”, we assert that that can’t be us, and that we in a sense live purely within the ruliad—which means that for us the Principle of Computational Equivalence holds, and we perceive only computation, not hypercomputation.
The is reminiscent of the question of the creator of God—“If God exists then who created God?”. Stephen’s rather unsatisfactory answer in the context of the hyperruliad existing is basically “we assert we are not in the hyperruliad because it’s not what we perceive” (presumably speaking on behalf of humankind).
Perhaps this problem can be solved if Aristotle’s challenge to the Pythagoreans is answered [metaphysicsaristotle]:
Those who say that existing things come from elements and that the first of existing things are the numbers, should have first distinguished the senses in which one thing comes from another, and then said in which sense number comes from its first principles.
Even granting abstractions without agency like numbers acting as a first cause, I will further show that observer-independent truth is impossible and so claims about external reality are void unless there is an explicit reconciliation. Consider the following objections a proponent of the theory might state:
O: hypergraphs don’t actually compose the ruliad
If not then there is a great disconnect between the models of the theory and the theory itself.
O: The ruliad itself is an abstraction—an abstraction of all transformations on hypergraphs
And so its instantiation is an article of faith. Then suppose Jonathan, a skeptic, disagrees on the grounds that the assumptions are not self-evident. Perhaps he is a determinist and rejects the claim we are “independent” from the universe. Using the model and language of Wolfram:
Stephen is an observer like us.
Stephen performs a lossy compression on a local space of rules adjacent to ours.
Stephen claims that $x$ is self-evident
Jonathan is an observer like us.
Jonathan performs a lossy compression on a local space of rules adjacent to ours.
[…]
O: Stephen and Jonathan can clear up their disagreement by proving x with a computational language
For problems in non-computational fields, Wolfram does not claim to have resolutions (but will soon!).
O: Grant that all the problems in our universe can be resolved by a computational language
In the model, truth is an emergent artifact of our nature as observers [Wolfram_truth]:
So what happens if one tries to add a statement that “isn’t true”? The basic answer is that it produces an “explosion” in which the entailment fabric can grow to encompass essentially any statement. From the point of view of underlying rules—or the ruliad—there’s really nothing wrong with this. But the issue is that it’s incompatible with an “observer like us”—or with any realistic idealization of a mathematician.
Because false statements and contradictory axioms exist in the ruliad, we must fall back and appeal to our nature, leaving the above disagreement up to some ill-defined consensus. Wolfram, by making the assumption of the “stability of discrete symbolic concepts of the kind of a computational language”, contradicts himself unless he can give an account for the faculty of abstract representation and its stability which doesn’t rely on the argument of usefulness from evolution (this would undermine its objectivity). But this is impossible because unlike in Tegmark’s universe, mathematical Truth is observer-dependent, and so perhaps disagreement—and moreover logic—is a kind of ambient babble.
In brief, we take the statement “the ruliad exists” as having a Truth value, but truth is according to our nature, and if it’s according to our nature, it’s not according to external reality and so does not have a Truth value—a contradiction.
O: Any human would find the ruliad reasonable given the common-sense assumptions
An appeal to common belief! External intelligence is also proposed in the model, so in the above example Jonathan can be replaced with an alien, or a rock. Is the ruliad no longer an objective concept if an alien mind could not have a conception of this ruliad? To go further, what if our nature as “observers like us” give rise to the abstraction of the ruliad?
O: Our nature as observers does give rise to the abstraction of the ruliad
Then the ruliad is about us. Then it can be true, false, or anything in between, if all is merely computation. At some point one would expect an extrapolation to reality independent of observers.
O: These are definitions of the observer which reproduce 20th century physics
Reproducing 20th century physics is reproducing models, not revealing the underlying structure of reality. And despite the purported potency of explaining the motion of bodies, Cartwright and Van Fraassen ask “What is it about explanation that guarantees truth?”. That being said, with the flexible combinatorial nature of the model it will likely be the case that many amazing phenomena will be able to be described by it.
O: The ruliad exists by formal necessity
But isn’t “formality”, a word also used in mathematics as well as philosophy, laden with an observers presuppositions? Gorard describes Definition 5 as coming from a formalism within mathematical logic, so it appears mathematical logic subsumes graph rewriting systems, and binary relations have an a priori existence undergirding them. Is there a causal chain connecting these concepts and how was this knowledge imprinted onto us particular observers? In Plato’s case, knowledge was recollection from the eternal soul. David Hilbert addressed mathematical theory as relations in 1899 [Blanchette_2007]:
It is surely obvious that every theory is only a scaffolding or schema of concepts together with their necessary relations to one another, and that the basic elements can be thought of in any way one likes. If in speaking of my points I think of some system of things, e.g., the system: love, law, chimney-sweep $\ldots$ and then assume all my axioms as relations between these things, then my propositions, e.g., Pythagoras’ theorem, are also valid for these things. In other words: any theory can always be applied to infinitely many systems of basic elements.
Wolfram has then applied this to his theory of relations between emes, the intangible atoms of computation. Relation comes from the Latin word referre which means “to carry back, notify, report”, which reminds us of colloquial usage when we “relate to” someone a message (it is also the root word for referee, referrant). I bring this up to frame the following question: how can one prove ontological priority of causal relations between emes over other relations (e.g. conferring a message between two humans, an emotional relation between two humans, a qualitative relation between colors)? Somehow, emes have privileged existence, otherwise they can be confidently relegated to the rest of the “coarse-grained” observations which comes from observers like us. I reject the argument that emes necessarily exist for there to be any kind of regularity or predictability in the universe due to insufficient evidence.
Figure 2: A causal graph laden with composite human-like ideas and a more abstract causal graph of unique identifiers. Which is which?
O: The ruliad exists in the same way 1+1=2 exists
[…]
Figure 3: Candidate rulial observers sampling the ruliad. As an observer one might reject one of these candidates, but hopefully the true observers will be known once the theory is complete.
An evolutionary biologist could claim that some humans may have been selected for based their ability to control external bodies in time and space, and to communicate. Then one can see how giving elevated—almost deified—status to the fundamental object (atoms) or to fundamental language (formal language) could be a byproduct of our evolution. Perhaps given this reductionist perspective a “great adjacency matrix in the sky” is a reasonable conclusion.
I will now briefly address some of Tegmark’s claims, and in doing so expound on some arguments already addressed above. Here are some key definitions given by Tegmark [Tegmark_book]:
Mathematical Universe Hypothesis (MUH): Our external physical reality is a mathematical structure.
Mathematical structure: Set of abstract entities with relations between them.
Equivalence: Two descriptions are equivalent if there’s a correspondence between them that preserves all relations.
Tegmark’s definition of equivalence only equates descriptions, and I posit that it would sound more absurd if instead the word objects was used. His language is vague by necessity—if someone told me the universe was abstract entities with relations between them I would assume this person was speaking about the spiritual realm. His intangible entities, like souls, also happen to be eternal. Furthermore his definition of equivalence is exactly that—it is not equality.
… If you can thus pair up every entity in our external physical reality with a corresponding one in a mathematical structure (“This electric-field strength here in physical space corresponds to this number in the mathematical structure,” for example), then our external physical reality meets the definition of being a mathematical structure—indeed, that same mathematical structure
Figure 4: There is a twoness inherent in both figures: in the top figure, the physical matter held in each hand is not identically equal and in the bottom figure—supposing a multiset—the cardinality is two. There is also a oneness in that relations between abstract components are preserved—the graphs are isomorphic. I believe both properties are real and eternal, but asserting oneness presupposed twoness and required subsequent computation.
Tegmark’s proof reduces to checking a lookup table of abstract models to physical measurements, which he expresses as a graph isomorphism problem (Fig. 4). This seems rather contrived and not my definition of identically equal: by using the phrase “these two things are identical” one already presupposes distinctness with the word two. Then one has to prove how his or her perception and process of enumeration failed to identify their oneness, rather than claiming they are identically equal because their components can be paired.
This aside, I assert that his method of proof places an impossible burden of computation on the observer and doesn’t escape the measurement problem. Tegmark’s theory radically diverges from Wolfram’s here in that Tegmark would need an observer that is not computationally bounded, a term used by Wolfram to describe the fact that despite knowing the rules or equations, we as humans can’t immediately predict certain futures. In spite of Tegmark’s view of an ultimately static universe, he still claims that the act of pairing up is required to prove his hypothesis. Then the question is: who is doing the pairing up, when, where, and how often? Graph isomorphism may be NP-complete and the number of vertices in this problem seems infinite. Tegmark acknowledges an internal reality subject to hallucinations, and how this is circumvented is unclear. Consider Dirac’s speculation [dirac]:
Dirac’s speculation: The constants of nature change over time $\approx$ the age of the universe.
This can still be consistent with a universe made of mathematics, but it confounds the claim that the math we have access to now, such as the Schrödinger equation, represents the immutable and eternal. There are always higher order corrections that can be woven into the narrative. To take a step back, pairing up requires the notion of binary relations and relation in mathematics is defined in the abstract. This, in and of itself, is not objectionable, but an account for the apparent and emergent faculty of humans to abstract is needed as abstract relations undergird the entire theory. So what does Tegmark mean by relation in his definition of mathematical structure? I would describe Max as an abstract entity with whom I have a relation. I would describe souls as abstract entities with relations between other souls, but although semantically correct, this is likely not within the scope of Tegmark’s definition.
O: The relations that describe Max are, in a first approximation, “human”, which is not the case for math abstractions.
Maybe Tegmark believes his human-independent description of the world, but as this theory comes from a human, I question the idea. In fact, abstraction is one of the most human things one can do. But I suppose a more apt description for Tegmark is a self-described “mathematical pattern in spacetime”. From Tegmark:
it must also be well defined according to nonhuman entities aliens or supercomputers, say—that lack any understanding of human concepts. Put differently, such a description must be expressible in a form that’s devoid of any human baggage like “particle”, “observation” or other English words.
or from Wolfram:
So we need to get rid of the “human-readable names” and just replace them with “lumps” $\ldots$
What baggage is in Tegmark’s definition of “baggage” and how does one conceive of “lumps” without instances? Because the computer is the tool du jour, I’ll make an analogy and say that this is like a class without an instance, or compile-time without run-time. Tegmark illustrates an example of non-humanness by comparing the description of a chess game using English language to one of algebraic chess notation, but it seems like an arbitrary choice to ascribe a more fundamental existence to the algebraic chess notation over, say, the narrative of the chess game.
Figure 5: A chess game and its compressed Forsyth–Edwards notation, converted to a byte string and then a binary number [Chojna_Warendorff_2023]. Wolfram and Tegmark think the binary notation is something more like the fabric of reality, whereas a chess player might want to look at the chess board and the players.
O: It is obvious that mathematical relations are significantly less human than, say, love between people
I personally cannot touch, feel, or perceive pure human love but I believe it exists in a similar platonic way that $1+1=2$ exists. It must be that Maxwell’s demon is picking and choosing what is allowed in our platonic realm.
O: Human love doesn’t predict motion of bodies in space and time
Prediction is a poor stand-in for truth.
O: Aliens and humans would have the same math, unlike platonic love
Wolram may disagree, asserting that math is intimately linked with our physical experience as observers. Furthermore, Tegmark’s supposed universal math has not been shown to be rigorous:
All of this begs the question: is it actually possible to find such a description of the external reality that involves no baggage? If so, such a description of objects in this external reality and the relations between them would have to be completely abstract, forcing any words or symbols to be mere labels with no preconceived meanings whatsoever. Instead, the only properties of these entities would be those embodied by the relations between them.
where the last sentence recapitulates Hilbert’s definition of theory as relations. So how do we come to terms with individuated and connected nothings giving rise to our experience?
O: Relations are emergent phenomena
Indeed, Tegmark asserts a kind of emergence:
The external physical reality is therefore more than the sum of its parts, in the sense that it can have many interesting properties while its parts have no intrinsic properties at all.
So after all the chapters describing the amazing reductionism in phenomena of physics (e.g. claiming that colors are actually electromagnetic spectra of Maxwell’s equations, everything is built of atoms), Tegmark asserts that properties can ex-nihilo emerge, but only from these particular “propertyless” entities. Would he grant such emergence for all other physics phenomena in opposition to their reductionist explanation? Along with his concept of perceptronium, perhaps he should introduce emergetronium as another distinct object, or else define emergence as a property, albeit latent, of mathematical structures.
O: These emergent properties are all artifacts of our consciousness
Tegmark believes consciousness is a “feeling”, but I would argue “feeling” takes place in consciousness and consciousness is needed to describe feeling in the first place. He continues:
So if the entities of this structure have no intrinsic properties, does the structure itself have any interesting properties (besides having eight elements)? In fact, it does: symmetries!
Now emergetronium must be invoked as there are not only relations, but by introducing symmetry Tegmark describes transformations on relations, like those on Wolfram’s hypergraphs.
Let us assume a person was born blind, deaf, and overall without senses. Perhaps they have sight, but exist in what we would describe as an empty space. What is this person’s understanding of symmetry, relations, feeling and enumeration? This person would only have a relation to “I am” which emerges from consciousness. Perhaps there’s some biological hard-wiring which would give them conceptions of symmetry, and this may confound the claim that math is non-human. Tegmark explores the theory of the observer through this lens of biological evolution:
We perceive that which is useful: Why do we perceive the world as stable and ourselves as local and unique? Here’s my guess: because it’s useful $\ldots$ Self-awareness would then be a side effect of this advanced information processing.
but doesn’t turn this reasoning about evolved usefulness on math itself. And in the reductionist evolutionary theory, self-awareness is a random mutation with no evolutionary advantage, because an organism could be unself-aware but still perform all the same motor movements and speech for procreation. By the same token, self-awareness could have evolved to be shared or infinitely-spanning. Tegmark makes the same circular guesses as Wolfram regarding the locality and uniqueness of the observer.
In a response to the question “Surely ‘stuff’ can’t be mathematical?”, Tegmark explains by making the analogy to a character in a video game who is not made of “stuff” and then writes [Tegmark]:
Sure, the computer in this example is itself made of stuff, but the feeling that the objects in the game were made of “stuff” was completely illusory and independent of the substrate out of which the computer was built.
I would like for him to think of an example that doesn’t actually have the “substrate out of which the computer is built”, rather than to have to explain this caveat. Then I would be satisfied, knowing that abstractions without agency can act as a first cause.
Wolfram clinically notes that “the only thing special about the universe to us is us ourselves” [Wolfram_toe] and I believe Tegmark, maintaining a great interest in ethics, would agree. Wolfram modifies the Cartesian partition (“God-World-I”) by reintroducing “I” as “observers” and attempts to meld them with “World”. Tegmark removes “I” and “God” and asserts that everything is in some sense “World”. I understand why Heisenberg warned that the Cartesian partition is dangerous. I respect both theories of everything, but find inconsistencies in the reconciliation of “I” the observer and the transcendent truths of mathematics—a cogent narrative has yet to be written which marries potential and instantiated truth. Aristotle thought that potentiality so understood is indefinable, claiming that the general idea can be grasped from a consideration of cases [aristotle]. Like Aristotle, I believe that actuality is to potential as
“what is awake is in relation to what is asleep, and what is seeing is in relation to what has its eyes closed but has sight, and what has been shaped out of the matter is in relation to the matter”
[…]